OpenStudy (anonymous):

PLEASE HELP! Need work shown please. 17 = p - 3 - 3p

OpenStudy (anonymous):

17 = p -3 -3p add 3 to both sides of the equation: 20 = p - 3p 20 = -2p divide both sides of the equation by -2 -10 = p p = -10

OpenStudy (anonymous):

THANK SO SO MUCH!!!!!!